# references.bib

### v5

@@ 1 Introduction
Current approaches to symbolic policy learning focus primarily on memory-less policies that have a fixed mapping from observations to control outputs [guo2024efficient, hein2018interpretable, nadizar2024naturally, landajuela2021discovering, gou2021knowledge, shimooka2000generating]. However, such policies may apply poorly to real-world settings, where volatility or partial observability complicate tasks. By adding memory to a control policy, past observations to improve policies in difficult settings [heess2015memory, khan2017memory]. This improvement has also been found by extending symbolic policies with memory [koza1992evolution, koza1999genetic, andersson1999reactive, raik1997evolving], where GP explicitly learns when to store and retrieve observations from memory.
@@ 2 Related Work
Before applying control policies to more realistic settings, the policies have to be able to handle partial observability and volatile environments. In [hein2018interpretable], the policy receives observations from past time points, which enables the policy to control a stochastic and partially observable task, although the policy itself does not contain explicit memory. This method requires storing a large number of observations, which does not scale well to longer trajectories or higher dimensions [khan2017memory], as the search space of GP becomes very large. A more general approach is to allow GP to learn to store computed values, which may be used for future computations [koza1999genetic, koza1992evolution, raik1997evolving, andersson1999reactive]. In [kelly2023discovering], control algorithms with explicit memory storage are evolved with genetic algorithms, and the resulting policies are able adapt to changes in the environment.
In contrast to these approaches, in this paper we will use GP to evolve the state equations of continuous-time latent variables that function as memory. Incoming observations are integrated into the dynamics of the latent variables, which offers a more general and scalable approach than using time-lagged observations [khan2017memory]. Instead of explicitly writing or reading values to memory, the latent variables are continuously updated by new observations and are added as an input to a readout that computes the control signal. This way, the policy can automatically select the relevant features to efficiently estimate hidden states, whereas using time-lagged observations would increase the number of irrelevant variables. The latent memory also improves estimation of variables from stochastic observations or adaptation environment settings. For clarity, in this paper the continuous-time memory-based policies are referred to as dynamic policies, and the memory-less policies as static policies.
The dynamic policies consist of multiple latent variables and a readout function, therefore multiple trees have to be optimised with GP simultaneously [langdon1998genetic]. In [cao2000evolutionary], GP evolves multiple trees that together represent a system of coupled differential equations. Although the trees could be learned independently [bongard2007automated], this setup is required when certain variables are not observed [cornforth2013inference]. This representation of multiple trees has also been utilised in collective decision making making [wen2016learning, brameier2001evolving] and multi-class classifiers [muni2004novel, ingalalli2014multi]. Based on these works, we extend our GP implementation to include tree based crossover and only apply crossover to pairs of trees from the same class.
@@ 3 Methods
The goal of this paper is to evolve dynamic symbolic policies that obtain high performance in control problems. We compare dynamic memory-based symbolic policies with static memory-less symbolic policies, which are explained in more detail. Afterwards, the specifics of the GP algorithm and how we extended GP to learn multiple trees efficiently are described. In the subsequent subsections the baselines and experiments we performed are detailed. For the implementation of GP, we utilised Kozax [^1] https://github.com/sdevries0/Kozax, a recently developed GP framework written in JAX. The code for reproducing the experiments and baselines is available at: https://github.com/sdevries0/DSP_paper.
@@ 3.1 Environment model
To evaluate a policy, a coupled system consisting of the policy and an environment is simulated, as described in Fig. 2a and Fig. 2b. In our paper, the policies are evaluated on dynamic tasks modelled as Markov decision processes. The environment is modelled by a state $\mathbf{x}\in\mathbb{R}^{N}$, following the dynamics of a stochastic differential equation
$d\mathbf{x}=f_{\textrm{env}}(\mathbf{x},\mathbf{u})dt+\mathbf{V}d\mathbf{w}$ (1)
@@
where $\mathbf{D}\in\mathbb{R}^{M\times N}$ is a readout matrix and $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$ is Gaussian observation noise. When relevant, the policy receives a target state $\mathbf{x}^{\ast}$ that should be achieved in the environment. The policies are evaluated on a batch of trajectories, with a batch size of 32. In every trajectory, the initial conditions are randomly sampled, and if applicable target states and environment parameters are sampled.
@@ 3.2.2 Dynamic policies
We extend this approach to dynamic symbolic policies (Fig. 2b), which implement memory through continuous-time dynamic latent variables. The latent variables process incoming observations, and the readout function maps the latent variables to a control signal. The dynamic policies contain expressions for the latent dimension and control dimension (Fig. 2c). The latent variables are governed by a system of ordinary differential equation, according to
@@
for $1\leq i\leq H$. Note that the dynamic policies also receive the previous control as an additional input. The readout functions of the dynamic policies are similar to Equation 3, but the input to $g_{j}$ is given by the latent variables $\mathbf{a}$ instead of the observations $\mathbf{y}$, which results in
@@ 3.4 Genetic programming
The symbolic policies are evolved via genetic programming. GP is a variant of evolutionary algorithms that focuses on learning computer programs [koza1994genetic]. A population of solutions is evolved to satisfy a certain goal or task through stochastic optimisation. In our setup, individuals are represented as tree-structured computational graphs, constructed from a predefined set of function nodes and leaf nodes. Function nodes cover mathematical operators and functions, while leaf nodes describe variables and constants. GP is largely defined by two hyperparameters. The first hyperparameter is the number of generations, which determines how many iterations of evolution are performed. The second hyperparameter is the population size, which defines the number of candidates that are evolved at every generation. See Algorithm A1 for an overview of the GP algorithm. The GP algorithm consists of initialisation, evaluation and reproduction steps, as described in more detail in the following sections.
@@ 3.4.4 Regularisation
Regularisation is necessary in GP to produce interpretable trees and improve the efficiency of the optimisation. To regularise the complexity of the reproduced trees, a punishment relative to the number of nodes in a tree is added to the fitness [luke2006comparison]. Additionally, the depth of the trees is limited to seven, and during initialisation the depth is set to four. To introduce more compact trees in the population, a small part of the population is simplified mathematically at every generation [hooper1996improving]. To improve diversity in the population during evolution, multiple subpopulations are evolved independently [fernandez2003empirical]. By configuring the hyperparameters in each subpopulation differently, exploration and exploitation is enforced explicitly [herrera2000gradual].
@@ 3.5 Baselines
Figure 3: Visualisations of the experiments. (a) In the harmonic oscillator, the position of the oscillator $x$ has to be moved to a target state $x^{*}$ by applying the control force $u$. (b) The acrobot consists of two links, which are each defined by the angles $\theta_{1}$ and $\theta_{2}$. To complete the swing up, the end of the second link has to be moved above the target height $x^{*}$ by applying control force $u$ on the second link. (c) The continuous stirred tank reactor is described by the concentration of reactant $c$, the temperature $T_{r}$ in the main reactor, and the temperature $T_{c}$ in the cooling jacket. The temperature in the reactor should be stabilised at a set point temperature $T_{r}^{*}$. The flow of coolant into the cooling jacket is controlled by $u$.
@@ 3.6 Environments
Our method and the baselines are evaluated on three different environments, as shown in Fig. 3. The policies are tested on a linear and non-linear classical control task and on a non-linear industrial problem. The chosen hyperparameter setting of the methods for each experiment are presented in Table 3 (Appendix 3.4). The choice of the function sets are described in the following subsections.
@@ 3.6.2 Acrobot swing up task
The acrobot is defined by two connected links, where the first link is connected to a fixed position (Fig. 3b). The goal is to swing the end point of the second link above a threshold, while minimising the applied force. The dynamics of the acrobot are non-linear, and to accomplish the swing up a non-linear controller is required. The policies solve the acrobot in experiments with noisy observations including and excluding the observations velocity, and a third experiment is conducted where control can be applied to both links individually. The observations of the acrobot include angles, therefore the sine and cosine functions are added to the function set of the GP algorithm, besides the arithmetic operators. The dynamics and the fitness function are described in Appendix C.
@@ 3.6.3 Continuous stirred tank reactor
The continuous stirred tank reactor (CSTR) describes a chemical process that is governed by non-linear dynamics [uppal1974dynamic]. The CSTR consists of a main tank that contains a mixture of chemicals and a cooling jacket around the main tank. A visualisation of the CSTR is presented in Fig. 3c. To optimise the chemical process inside the main tank, the temperature has to be stabilised around a setpoint temperature. The temperature in the cooling jacket can be influenced directly by altering the input flow of coolant, which in turn affects the temperature in the main tank. in our experiment, the parameters of the CSTR are varied in every trial (Table III). The policy receives noisy observations of the temperature in the reactor and cooling jacket, but the concentration is not observed. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid expressions to deal with the values of the temperatures, which typically range between 200 and 600. The dynamics and the values of the environment’s parameters are detailed in Appendix D, as well as the fitness function.
@@ 4 Results
We present a GP method for finding high-performing dynamic symbolic policies for solving control problems. The performance of a policy is tested by controlling three common control benchmark systems (as explained in detail in Section 3.6). We investigate whether GP evolves successful control policies for linear and non-linear environments, different observability levels, as well as changing environment conditions and multiple control dimensions. We compare the resulting dynamic symbolic policies to static symbolic policies, optimal control (where possible), random search (RS), and to a reference black-box model (NDEs).
Figure 4: Genetic programming evolves high-performing dynamic symbolic policies. (a) Evolution results of the experiment on the stochastic harmonic oscillator (SHO) with observation noise. The left plot shows the best fitness at every generation during evolution, averaged over 20 independent runs. The methods included are genetic programming for static (GP-S) and dynamic (GP-D) policies, random search (RS), neural differential equation (NDE) and the analytical optimal linear quadratic Gaussian (LQG). The right plot presents the validation fitness obtained on unseen conditions with the best policy evolved in every evolutionary run. (b) Evolution results of the experiment on the acrobot with noisy observations. The LQG is excluded because it cannot be computed given non-linear dynamics. (c) Evolution results for the experiment in which the state of the SHO is partially observable. (d) Evolution results of the experiment on the acrobot under partial state observability. (e) Evolution results of the experiment on the SHO with varying parameters. (f) Evolution results of the experiment on the acrobot with two control inputs.
To study whether GP can successfully evolve dynamic symbolic policies, we first consider the SHO and acrobot environments with noisy full state observability (Fig. 4a,b). In the SHO, we consider the linear quadratic Gaussian (LQG) controller as the upper bound fitness (black curves). The fitness of random search improves during optimisation (purple curves), but does not converge near the upper bound nor the other methods. Random search does occasionally discover a policy that competes with the evolutionary approaches, but not consistently (purple dots). Moving forward, we will no longer consider random search as a baseline, but the results were similar across all tasks. NDEs converge to a high fitness quickly and consistently, approximating the upper bound fitness in the SHO (green curves and dots).
To validate that symbolic policies reach similar performance as the NDE, we first optimised static symbolic policies (blue curves). The initial fitness of GP-S begins close to the upper bound fitness in the SHO, which indicates that well-performing static policies exist in the initial population, which is a result of the small search space of GP-S. However, from there, GP is not able to evolve policies reaching the optimal fitness, indicating that static policies are not complex enough to match the optimal controller. Similarly, in the acrobot swing-up task, GP-S begins at a high fitness, but does not improve much during evolution. In both tasks, GP-S converges significantly higher than RS, but is outperformed by the NDE in terms of fitness.
We next evolved dynamic symbolic policies (orange curves). In both tasks, the fitness of GP-D begins significantly lower than GP-S, due to the larger search space. However, GP-D quickly converges to a higher fitness than GP-S, approximating the upper bound fitness in the SHO. This indicates that the internal dynamics are able to correct for the noisy observations. Both the GP-S and GP-D policies are highly consistent across trials, with GP-D producing slightly better policies (blue and orange dots).
Overall, these experiments demonstrates that static and dynamic symbolic policies are effectively optimised with GP, especially compared to random search. High-performing linear and non-linear policies are evolved that compete with the black-box NDE, although the NDE is more consistent.
The previous experiments demonstrated that static and dynamic symbolic policies are successfully evolved for linear and non-linear settings. However, these experiments assumed the full state is observed, which may not be representative for real-life applications. We hypothesised that the memory in dynamic policies is particularly beneficial when confronted with partial state observations. To test this hypothesis, symbolic policies were evolved for the SHO and acrobot environments under partial state observability, where the velocities are not observed (Fig. 4c,d).
The SHO remains linear, therefore the LQG controller again represents the upper bound fitness (black curves). The NDE converges to a slightly lower fitness (green curves) compared to the experiments with full state observability. However, the NDE still consistently produced policies close to the upper bound in terms of performance, with a few policies that perform worse on the acrobot (green dots).
As hypothesised, GP-S converges to a substantially lower fitness than under full state observability (blue curves). Every static symbolic policy fails to satisfy the task when the velocity is not observed (blue dots). Differently from GP-S, high-performing dynamic symbolic policies are still successfully evolved (orange curves). GP-D converges to a similar fitness as the NDE in both environments. Additionally, the validation fitness is consistent, and some policies even perform better than the NDE policies (orange dots).
@@ 4.2 Dynamic symbolic policies remain robust under partial state observability
Table 1: The fitness of the best policies on validation trajectories for every experiment. The complexity of the best policies is also displayed as the size. The methods include random search (RS), neural differential equation (NDE) and genetic programming for static (GP-S) and dynamic (GP-D) policies. The bold fitness indicates that a method obtained the best fitness in an experiment. The complexity is computed by summing the number of mathematical operations, variables and constants in the policies.
@@
SHO Partial state observability - -72.76 195 -903.44 5 -63.35 21
SHO Varying parameters - -47.76 215 -110.25 13 -45.56 25
@@
Acrobot Partial state observability - -101.47 191 -250.0 1 -110.23 12
Acrobot Two control inputs - -97.02 272 -103.43 12 -100.69 18
CSTR Standard - -1017.72 215 -1059.31 9 -845.33 27

### v6

@@ 1 Introduction
Current approaches to symbolic policy learning focus primarily on memory-less policies that have a fixed mapping from observations to control outputs [guo2024efficient, hein2018interpretable, nadizar2024naturally, landajuela2021discovering, gou2021knowledge, shimooka2000generating]. However, such policies may apply poorly to real-world settings, where volatility or partial observability complicate tasks. By adding memory to a control policy, past observations to improve policies in difficult settings [heess2015memory, khan2017memory]. This improvement has also been found by extending symbolic policies with memory [koza1992evolution, koza1999genetic, andersson1999reactive, raik1997evolving], where GP explicitly learns to store and retrieve observations from memory.
@@ 2 Related Work
In real-world settings, control policies should be able to handle partial observability and volatility in the environment. In [hein2018interpretable, hein2021trustworthy], the policy receives observations from past time points, which enables the policy to control a stochastic and partially observable task, although the policy itself does not contain explicit memory. A method based on lagged observations requires storing a large number of data points, which does not remain robust when observations are noisy or in case of long trajectories [khan2017memory, lin1993reinforcement, mccallum1996hidden], as this leads to a large increase of the search space in GP. A more general approach is to allow GP to learn to store computed values, which may be used for future computations [koza1999genetic, koza1992evolution, raik1997evolving, andersson1999reactive]. In [kelly2023discovering], control systems with explicit memory storage are evolved with genetic algorithms, and the resulting policies are able adapt to changes in the environment.
Contrary to these implementations of memory in control policies, in this paper we will use GP to evolve the state equations of continuous-time latent variables that function as memory. Similar approaches apply GP to optimise the structure of recurrent neural networks with internal memory [turner2017recurrent, esparcia1997evolving]. However in these works, the evolved networks operate in discrete time, individual nodes in the networks have values that are updated at each time step and the networks do not consist of symbolic equations. In our approach, incoming observations are continuously integrated into the dynamics of the latent variables and function as (recurrent) inputs for the functions in the symbolic control policy at any point in time. For clarity, in this paper the continuous-time memory-based policies are referred to as dynamic policies, and the memory-less policies as static policies.
The dynamic policies consist of multiple latent variables and readout functions, therefore multiple trees have to be optimised with GP simultaneously [langdon1998genetic]. In [cao2000evolutionary], GP evolves multiple trees that together represent a system of coupled differential equations. Although the trees could be learned independently [bongard2007automated], this setup is required when certain variables are not observed [cornforth2013inference]. This representation of multiple trees has also been utilised in collective decision making making [wen2016learning, brameier2001evolving] and multi-class classifiers [muni2004novel, ingalalli2014multi]. Based on these works, we extend our GP implementation to include tree based crossover and only apply crossover to pairs of trees from the same class.
@@ 3 Methods
The goal of this paper is to evolve dynamic symbolic policies that obtain high performance in control problems. We compare dynamic memory-based symbolic policies with static memory-less symbolic policies, which are explained in more detail. Afterwards, the specifics of the GP algorithm and how we extended GP to learn multiple trees efficiently are described. In the subsequent subsections the baselines and experiments we performed are detailed. For the implementation of GP, we utilised an initial version of Kozax [de2025kozax], a recently developed GP framework written in JAX. The code for reproducing the experiments and baselines is available at: https://github.com/sdevries0/DSP_paper.
@@ 3.1 Environment model
To evaluate a policy, a coupled system consisting of the policy and an environment is simulated, as described in Fig. 2a and Fig. 2b. In our paper, the policies are evaluated on dynamic tasks modelled as Markov decision processes. Throughout this paper, we will denote $x(t)$ as $x$ without explicit time dependence for notational convenience. The environment is modelled by a state $\mathbf{x}\in\mathbb{R}^{N}$, following the dynamics of a stochastic differential equation
[…]
@@
where $\mathbf{D}\in\mathbb{R}^{M\times N}$ is a readout matrix and $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$ is Gaussian observation noise, where $\mathbf{\Sigma}=\sigma\mathbf{I}$. When relevant, the policy receives a target state $\mathbf{x}^{\ast}$ that should be achieved in the environment. The policies are evaluated on a batch of trajectories, with a batch size of 32. In every trajectory, the initial conditions are randomly sampled, and if applicable target states and environment parameters are sampled.
@@ 3.2.2 Dynamic policies
We extend this approach to dynamic symbolic policies (Fig. 2b), which implement memory through continuous-time dynamic latent variables. The latent variables process incoming observations, and the readout function maps the latent variables to a control signal. The dynamic policies contain expressions for the latent dimension and control dimension (Fig. 2c). The latent variables are governed by a system of ordinary differential equation according to
@@
[…]
@@ 3.4 Genetic programming
The symbolic policies are evolved via genetic programming. GP is a variant of evolutionary algorithms that focuses on learning computer programs [koza1994genetic]. A population of solutions is evolved to satisfy a certain goal or task through stochastic optimisation. In our setup, individuals are represented as tree-structured computational graphs, constructed from a predefined set of function nodes and leaf nodes. Function nodes cover mathematical operators and functions, while leaf nodes describe variables and constants. Two important hyperparameters in GP are the number of generations, which determines how many iterations of evolution are performed, and the population size, which defines the number of candidates that are evolved at every generation. These hyperparameters should be balanced to reach convergence while remaining computationally efficient. See Algorithm A1 for an overview of the GP algorithm. The GP algorithm consists of initialisation, evaluation and reproduction steps, as described in more detail in the following sections.
@@ 3.4.4 Regularisation
Regularisation is necessary in GP to produce interpretable trees and improve the efficiency of the optimisation. To regularise the complexity of the reproduced trees, a punishment relative to the number of nodes in a tree is added to the fitness [luke2006comparison]. Additionally, the depth of the trees is limited to seven, and during initialisation the depth is set to four. To introduce more compact trees in the population, a small part of the population is simplified mathematically at every generation [hooper1996improving]. The equations are simplified using SymPy [meurer2017sympy], after which they are restructured into the tree representation. To improve diversity in the population during evolution, multiple subpopulations are evolved independently [fernandez2003empirical]. Each subpopulation is evolved with different selection pressure, probability of a tree to be changed during evolution and probabilities for crossover, mutation, random sampling and simplification. This way, subpopulations with high crossover and simplification probability focus more on exploitation, while exploration is enforced for subpopulations with high mutation and random sampling rates, as well as higher selection pressure and probability to evolve a tree in a candidate. Configuring subpopulations with different hyperparameters to balance exploration and exploitation is beneficial [herrera2000gradual], especially combined with migration between the subpopulations.
@@ 3.6 Environments
Our method and the baselines are evaluated on three different environments. The policies are tested on a linear and non-linear classical control task and on a non-linear industrial problem. The number of generations, population size and latent state size used in each experiment are presented in Appendix Table 3. The choice for the population size and number of generations were chosen to balance efficiency and consistent convergence empirically. To ensure that the policies have enough computational complexity, number of latent states in the dynamic symbolic policies was set to two in every experiment that requires one control input, and set to three in the experiment with multiple control inputs. The choice of the function sets are described in the following subsections.
@@ 3.6.1 Stochastic harmonic oscillator
Figure 3: Genetic programming evolves high-performing dynamic symbolic policies for the stochastic harmonic oscillator. (a) In the stochastic harmonic oscillator (SHO), the position of the oscillator $x$ has to be moved to a target state $x^{*}$ by applying the control force $u$. (b) Evolution results of the experiment on the SHO with observation noise. The left plot shows the best fitness at every generation during evolution, averaged over 20 independent runs. The methods included are genetic programming for static (GP-S) and dynamic (GP-D) policies, random search (RS), neural differential equation (NDE) and the analytical optimal linear quadratic Gaussian (LQG). The right plot presents the validation fitness obtained on unseen conditions with the best policy evolved in every evolutionary run. (c) Evolution results for the experiment in which the state of the SHO is partially observable. (d) Evolution results of the experiment on the SHO with varying environment parameters.
@@ 3.6.2 Acrobot swing up task
The acrobot is defined by two connected links, where the first link is connected to a fixed position (Fig. 5a). The goal is to swing the end point of the second link above a threshold, while minimising the applied force. The dynamics of the acrobot are non-linear, and to accomplish the swing up a non-linear controller is required. The policies solve the acrobot in experiments with noisy observations including and excluding the observations velocity, and a third experiment is conducted where control can be applied to both links individually. The observations of the acrobot include angles, therefore the sine and cosine functions are added to the function set of the GP algorithm, besides the arithmetic operators. The dynamics and the fitness function are described in Appendix C.
@@ 3.6.3 Continuous stirred tank reactor
The continuous stirred tank reactor (CSTR) describes a chemical process that is governed by non-linear dynamics [uppal1974dynamic]. The CSTR consists of a main tank that contains a mixture of chemicals and a cooling jacket around the main tank. A visualisation of the CSTR is presented in Fig. 6a. To optimise the chemical process inside the main tank, the temperature has to be stabilised around a setpoint temperature. The temperature in the cooling jacket can be influenced directly by altering the input flow of coolant, which in turn affects the temperature in the main tank. in our experiment, the parameters of the CSTR are varied in every trial (Table III). The policy receives noisy observations of the temperature in the reactor and cooling jacket, but the concentration is not observed. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid expressions to deal with the values of the temperatures, which typically range between 200 and 600. The dynamics and the values of the environment’s parameters are detailed in Appendix D, as well as the fitness function.
@@ 4 Results
We present a GP method for finding high-performing dynamic symbolic policies for solving control problems. The performance of a policy is tested by controlling three common control benchmark systems (as explained in detail in Section 3.6). We investigated whether GP evolves successful control policies for linear and non-linear environments, different observability levels, as well as changing environment conditions and multiple control dimensions. We compared the resulting dynamic symbolic policies to static symbolic policies, optimal control (where possible), random search (RS), and to a reference black-box model (NDEs).
@@
Table 1: Fitness of the best policies on validation trajectories for every experiment. The complexity of the best policies is also displayed as the size. The methods include random search (RS), neural differential equation (NDE) and genetic programming for static (GP-S) and dynamic (GP-D) policies. The bold fitness indicates that a method obtained the best fitness in an experiment. The complexity is computed by summing the number of mathematical operations, variables and constants in the policies.
@@
SHO Partial state observability - - -72.76 195 -903.44 5 -63.35 21
SHO Varying parameters - - -47.76 215 -110.25 13 -45.56 25
@@
Acrobot Partial state observability - - -101.47 191 -250.0 1 -110.23 12
Acrobot Two control inputs - - -97.02 272 -103.43 12 -100.69 18
CSTR Standard - - -1017.72 215 -1059.31 9 -845.33 27
To study whether GP can successfully evolve dynamic symbolic policies, we first consider the SHO and acrobot environments with noisy full state observability (Figs. 3b and 5b). In the SHO, we define the linear quadratic Gaussian (LQG) controller as the upper bound fitness (horizontal black lines). The fitness of random search improved during optimisation (purple curves), but did not converge near the upper bound nor the other methods. Random search did occasionally discover a policy that competes with the evolutionary approaches, but not consistently (purple dots). Moving forward, we will no longer consider random search as a baseline, but the results were similar across all tasks. NDEs converged to a high fitness quickly and consistently, approximating the upper bound fitness in the SHO (green curves and dots).
To validate that symbolic policies reach similar performance as the NDE, we first optimised static symbolic policies (blue curves). The initial fitness of GP-S started close to the upper bound fitness in the SHO, which indicates that well-performing static policies exist in the initial population, which is a result of the small search space of GP-S. However, from there, GP was not able to evolve policies reaching the optimal fitness, indicating that static policies are not complex enough to match the optimal controller. Similarly, in the acrobot swing-up task, GP-S began at a high fitness, but did not improve much during evolution. In both tasks, GP-S converged significantly higher than RS, but was outperformed by the NDE in terms of fitness.
We next evolved dynamic symbolic policies (orange curves). In both tasks, the fitness of GP-D began significantly lower than GP-S, due to the larger search space. However, GP-D quickly converged to a higher fitness than GP-S, approximating the upper bound fitness in the SHO. This indicates that the internal dynamics are able to correct for the noisy observations. Both the GP-S and GP-D policies are highly consistent across trials, with GP-D producing slightly better policies (blue and orange dots).
Overall, these experiments demonstrate that static and dynamic symbolic policies are effectively optimised with GP, especially compared to random search. High-performing linear and non-linear policies were evolved that compete with the black-box NDE, although the NDE is more consistent.
The previous experiments demonstrated that static and dynamic symbolic policies are successfully evolved for linear and non-linear settings. However, these experiments assumed the full state is observed, which may not be representative for real-life applications. We hypothesised that the memory in dynamic policies is particularly beneficial when confronted with partial state observations. To test this hypothesis, symbolic policies were evolved for the SHO and acrobot environments under partial state observability, where the velocities are not observed (Figs. 3c and 5c).
The SHO remains linear, therefore the LQG controller again represents the upper bound fitness (horizontal black lines). The NDE converged to a slightly lower fitness (green curves) compared to the experiments with full state observability. However, the NDE still consistently produced policies close to the upper bound in terms of performance, with a few policies that perform worse on the acrobot (green dots).
As hypothesised, GP-S converged to a substantially lower fitness than under full state observability (blue curves). Every static symbolic policy fails to satisfy the task when the velocity is not observed (blue dots). Differently from GP-S, high-performing dynamic symbolic policies were still successfully evolved (orange curves). GP-D converged to a similar fitness as the NDE in both environments. Additionally, the validation fitness is consistent, and some policies even perform better than the NDE policies (orange dots).
Figure 4: Dynamic policies remain more robust than time-lagged policies under higher levels of noise. Policies without memory (GP-S), with latent memory (GP-D) and access to time-lagged variables (GP-L) were evolved on the stochastic harmonic oscillator (SHO) with partial state observability and varying levels of observation noise $\sigma$. The average validation fitness on unseen conditions is shown for the policies evolved with 20 different seeds. The time-lagged policies received observations from the last three time steps as input.
Figure 5: Genetic programming evolves high-performing dynamic symbolic policies for the acrobot swing-up task. (a) The acrobot consists of two links, which are each defined by the angles $\theta_{1}$ and $\theta_{2}$. To complete the swing up, the end of the second link has to be moved above the target height $x^{*}$ by applying control force $u$ on the second link. (b) Evolution results of the experiment in the acrobot environment with observation noise. The left plot shows the best fitness at every generation during evolution, averaged over 20 independent runs. The methods included are genetic programming for static (GP-S) and dynamic (GP-D) policies, random search (RS) and neural differential equation (NDE). The right plot presents the validation fitness obtained on unseen conditions with the best policy evolved in every evolutionary run. (c) Evolution results of the experiment on the acrobot under partial state observability. (d) Evolution results of the experiment on the acrobot with two control inputs.
Naively applying static policies to partially observable environments is expected to perform poorly. In practice, researchers would provide time-lagged observations to improve control policies [hein2018interpretable, mnih2013playing, uehara2022provably]. Lagged variables in principle allow programs to infer more information and deal with e.g. partial observability by estimating missing variables. However as the level of noise increases, more history of observations are required for accurate estimation. Consequently, the search space of GP increases with more observations and performance deteriorates with limited history length. The latent dynamic memory proposed in this work does not suffer from this limitation, as observations are continuously integrated. To demonstrate this, Fig. 4 shows the comparison between GP-S, GP-D and policies that receive time-lagged observations (GP-L, pink dots) on the SHO with partial state observability and increasing $\sigma$ (Eq. 2), representing more noise on the observation of the position. GP-L received the three most recent observations of the position at every time step. With these time-lagged observations of the position, the velocity can be estimated and used for accurate control. Fig. 4 shows that without observation noise, GP-L (pink dots) is competitive with GP-D (orange dots) and outperforms GP-S (blue dots). However as the observation noise increased, it became more difficult to estimate the velocity and the performance of GP-L deteriorates, moving closer to GP-S in terms of performance. GP-D remained constant in performance when the observation noise increased, demonstrating more reliable estimation capabilities through continuous integration of observations. The best policies of GP-L, GP-S and GP-D are presented in Appendix Table 6 for every noise level. Increasing the number of time-lagged observations could improve the estimation given high noise levels, but also scales the size of the search space. The dynamic policies generalise the time-lagged observations, as the latent variables can integrate an arbitrary length of time steps.